OpenStudy (anonymous):

y varies inversely as x. When y = 3, x = 12. Determine y when x = 2.

OpenStudy (igreen):

3 / 12 = 0.25 Therefore, when x = 2, y = 2 * 0.25 = 1/2.

OpenStudy (anonymous):

60/2=30 Remeber 60/x=y

OpenStudy (anonymous):

What are the answer choices?

OpenStudy (anonymous):

http://www.mathwords.com/i/inverse_variation.htm

OpenStudy (anonymous):

a) 8 b) 11 c) 18 d) 24